Project

General

Profile

2017-02-28 physics-shifter

Tasks:
  • Before Thursday - recovery from quench
  • Wednesday-Thursday 23rd Feb - SSU and SSD only, powered in flip mode; momentum scan for magnet alignment
  • Thursday 24th Feb - Sunday 26th Feb - magnets ramped to 2016/05 1
    • momentum scan then emittance scan
    • diffuser 4 stuck on highest emittance setting
  • Sunday 26th Feb - Tuesday 28th Feb - ISIS down for failure of dipole 9
  • Emittance Reduction

Magnet alignment

  • Noted issues with track extrapolation from detectors
  • For SSs only data, SSD was marked as solenoid mode in the reconstruction software; resolved, awaiting reconstruction reprocessing
  • For 2016/05 1 data, SSD Trim coils were marked as 0 current; resolved, awaiting reconstruction reprocessing

Emittance Reduction

  • Note that DiffMask8-3-140+M3-TSpecial Run is misaligned
Tag Runs lmc1234 sum TOF1 Triggers TOF2 Triggers time taken number in muon peak number at tof2 number passing all cuts number passing all cuts / p in tku [MeV/c^-1] number passing all cuts / tof01 [ns^-1]
DiffMask8-3-140+M3-TSpecial Run [9209] 610282 150483 46090 1 hrs 58 mins 49146 18522 13312 1438.62 25815.0
10-140+M3-Test3 [9203] 326423 245332 37004 0 hrs 53 mins 64362 12404 9521 1099.26 20965.0

Sample output

bunch_x_y.png - DiffMask8-3-140+M3-TSpecial Run (23.6 KB) Rogers, Chris, 28 February 2017 14:22

bunch_x_px.png - DiffMask8-3-140+M3-TSpecial Run (23.9 KB) Rogers, Chris, 28 February 2017 14:22

bunch_y_py.png - DiffMask8-3-140+M3-TSpecial Run (24.3 KB) Rogers, Chris, 28 February 2017 14:22